Michael mows lawns for his summer job. He only averages three lawns a week for the first right weeks, but then averages six lawns a week for the last four. What is the average number of lawns he mowed per week during the entire summer?
Answer by jorel1380(3719)

8x3+6x4/12=24+24/12=48/12=4 lawns per week average..